OpenStudy (anonymous):

Evaluate: 10 − 64 ÷ 8 + 22

OpenStudy (anonymous):

division first, then do the operations from left to right.

Parth (parthkohli):

P E M D A S

Parth (parthkohli):

^ order of operations, as it is known

Parth (parthkohli):

\( \color{Black}{\Rightarrow 10 - [(64 \div 8) + 22] }\)

OpenStudy (zarkon):

I wouldn't put the 22 in the brackets

OpenStudy (anonymous):

I Keep getting 24 but thats not correct

OpenStudy (anonymous):

10-30= -20

OpenStudy (anonymous):

start with 64 divided by 8... which is 8 . then bring every thing else down 10-8+22 then subtract 10-8 which is 2 then add 2+22 which is 24... idk why you say 24 is incorrect....

OpenStudy (anonymous):

oh im sorry the 2nd 2 is squared

Parth (parthkohli):

If squared, then do the exponent first.

OpenStudy (anonymous):

64 divide into 8 8+22=30-10 =-20

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

It's 6

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

the 2 is squared do exponet first you get 4 then follow pemdas

OpenStudy (anonymous):

instead of 22 the 2nd 2 is squared in the problem